OpenStudy (anonymous):

Lauren, Marie, and Clarissa were doing sit-ups in their physical education class. They counted how many sit-ups they could do in a minute. Lauren did twice as many sit-ups as Marie. Marie did five fewer sit-ups than Clarissa. If Clarissa did 33 sit-ups in a minute, how many sit-ups did Lauren do?

OpenStudy (studyhard):

Clarissa: 33 situps. Marie(5 fewer): 33-5=28 situps. Lauren did twice as many as Marie=28*2=56. Therefore, Lauren did 56 situps.